the angles of the triangle are abc. angle b is twice as angle a. angle c is 12 times as angle b. what are the measures of all angles?
angle abc = 180

Answer by edjones(7569)

b=2a c=12b=24a
a+2a+24a=27a
27a=180
a=20/3=6 2/3
b=40/3=13 1/3
c=480/3=160
Ed